Thirsty Cactus Corp. just paid a dividend of $1.45 per share. The dividends are expected to grow at 40 percent for the next 7 years and then level off to a 6 percent growth rate indefinitely.

   

Required :
If the required return is 12 percent, what is the price of the stock today?

$3.26

$149.47

$122.15

$152.46

$146.48
Answer #1

D1=(1.45*1.4)=2.03

D2=(2.03*1.4)=2.842

D3=(2.842*1.4)=3.9788

D4=(3.9788*1.4)=5.57032

D5=(5.57032*1.4)=7.798448

D6=(7.798448*1.4)=10.9178272

D7=(10.9178272*1.4)=15.28495808

Value after year 7=(D7*Growth Rate)/(Required rate-Growth Rate)

=(15.28495808*1.06)/(0.12-0.06)

=$270.0342594

Hence current price=Future dividends and value*Present value of discounting factor(rate%,time period)

=2.03/1.12+2.842/1.12^2+3.9788/1.12^3+5.57032/1.12^4+7.798448/1.12^5+10.9178272/1.12^6+15.28495808/1.12^7+$270.0342594/1.12^7

=$149.47(Approx).